OpenStudy (oswinoswald):

Advanced algebra help? Could also be considered finance...? Can someone give me an example for how to calculate the monthly payment on an ARM?

OpenStudy (homework_help):

http://www.bankrate.com/calculators/mortgages/adjustable-rate-mortgage-arm-calculator.aspx This is a site that will help you calculate monthly payment if you have the numbers to put in. Are you looking to know how to do it without any help and just by yourself with pen and paper? Or would this site help?

OpenStudy (oswinoswald):

I am looking for how to do it with pen and paper. I've done it before for this maths class, but I forgot how.

OpenStudy (mathmate):

Here's another site that gives you hands-on example. http://www.investinganswers.com/financial-dictionary/real-estate/adjustable-rate-mortgage-arm-41

OpenStudy (oswinoswald):

Thank you, @mathmate, but this still doesn't help. I need step-by-step directions on how to do it. I came on OS because I couldn't find what I needed online.

OpenStudy (mathmate):

The best way is to set up an example and work on it with pen and paper. Say you have a 100,000 loan. What kind of ARM do you want to try? ARM which varies every month, say prime + 3 %, ok?

OpenStudy (oswinoswald):

Well, the problem that I have to solve is a 30 yr 4/2 ARM with a 4/7 cap... So something like that?

OpenStudy (rperry309):

cymath.com

OpenStudy (mathmate):

So it's a 30 yr. mortgage, 4/2 ARM means that the first four years have a fixed (determined) interest rate, and every 2 years thereafter, the rate is adjusted. A 4/7 cap means that the first interest adjustment, the maximum change (+ or -) cannot exceed 4% (initial cap), and thereafter, a maximum change of 7% (life-time cap) every two years. So far so good?

OpenStudy (rperry309):

http://www.cymath.com/

OpenStudy (rperry309):

https://www.wolframalpha.com/

OpenStudy (rperry309):

those help if you can put the problem in number then those websites will do the rest of the work

OpenStudy (oswinoswald):

Looks good, I understand that stuff. I'm just confused as to the difference in the process to figuring out the monthly payments on a fixed rate mortgage. @mathmate

OpenStudy (oswinoswald):

@rperry309 thanks, but if you'd read previous comments, you'd see that's not what I'm looking for.

OpenStudy (rperry309):

nvm ion know that stuff thats big kid stuff

OpenStudy (oswinoswald):

So, with all due respect, please go away if you're not going to read through before you try to help. @rperry309

OpenStudy (mathmate):

@OswinOswald Once you have the mortgage defined, you can proceed using very simple calculations. It's a long calculation by pen and paper because of the variation of rates, you cannot figure out with a single formula. You need to figure out monthly (at least yearly) over thirty years. The best compromise is the work out the problem using Excel, and change the rates as they come. In your case, there will be 13 changes in rate (4 initial + 13 changes make 30 years).

OpenStudy (mathmate):

If you wish, we can do a 3/1 ARM with a 2/2 cap as an example.

OpenStudy (oswinoswald):

Sure.

OpenStudy (mathmate):

So let's say we have a 100,000 mortgage, with initial rate 4% over 3 years, followed by 5,7%, i.e. 4,4,4,5,7% over 5 years. So far so good?

OpenStudy (oswinoswald):

Okay...

OpenStudy (oswinoswald):

So, I guess my question is how do you figure out the initial period (or any single period), because with the regular monthly payment equation, you're figuring for paying off the entire thing in a set period of time. With ARM, you don't pay the whole thing in the first period, so how do you account for that? If this doesn't make sense, just ignore it and proceed haha.

OpenStudy (mathmate):

I am not sure of the answer, and I am in the process of double checking. I believe we calculate the payment based on the current rate for the remaining term...until the rate changes. But I will double check using the example in http://www.quickenloans.com/blog/adjustable-rate-mortgage-terms-arm-knowledge Will take a few minutes, but will be back.

OpenStudy (oswinoswald):

I'm gonna close this question while you're doing that, so I can ask another one. Please let me know if you figure it out.

OpenStudy (mathmate):

Sure!

OpenStudy (oswinoswald):

Here's the actual question: Jackie obtains a 30-year 4/2 ARM at 5% with a 4/7 cap structure in the amount of $313,500. What is the monthly payment during the initial period?

OpenStudy (oswinoswald):

I can find the answer on previous OpenStudy posts, but that doesn't really help for when I get more questions like this...

OpenStudy (oswinoswald):

@mathmate ?

OpenStudy (mathmate):

It is much easier for the initial period. The monthly payment is exactly the same as a regular mortgage, with the same principle, interest equals the initial rate, and the period is the amortizing period. So 30-year 4/2 ARM at 5% with a 4/7 cap structure in the amount of $313,500 For the initial period, assumed compounded monthly, Principal, P=313500, Amortized over 30 years, n = 360 months (periods) initial interest rate, i = 5% p.a. = 0.05/12 per month. initial period = 4 years (does not really matter). Monthly payment is calculated by the usual formula \(\large A =P\frac{(1+i)^n(i)}{(1+i)^n-1}\) \(\large ~~ =313500\frac{(1+0.05/12)^{360}(0.05/12)}{(1+0.05/12)^{360}-1}\) \(\Large ~~= 1682.9358\) per month. Hope that helps!

OpenStudy (mathmate):

By the way, if you searched OS, and found one post by Anonymous, the answer given $1832 is incorrect! Use the equation I gave above.